given the equal of the wave y(x, t)= 1.0mcos(2πm^-1 ×-π/3s t)


1
Expert's answer
2021-01-30T08:17:01-0500

a)


"\lambda=\dfrac{2\pi}{k}=\dfrac{2\pi}{2\pi\ m^{-1}}=1.0\ m."

b)


"f=\dfrac{\omega}{2\pi}=\dfrac{\dfrac{\pi}{3} \dfrac{rad}{s}}{2\pi}=0.166\ Hz."

c)


"\omega=\dfrac{\pi}{3}\ \dfrac{rad}{s}."

d)


"T=\dfrac{1}{f}=\dfrac{1}{0.166\ Hz}=6.0\ s."

e)


"v=f\lambda=0.166\ Hz\cdot1.0\ m=0.166\ \dfrac{m}{s}."